OpenStudy (anonymous):

What is the value of the expression? -(2/5)-(3/7)*(7/10) A. -7/10 B. -1/3 C. 1/3 D. 7/10

OpenStudy (anonymous):

-2/2 = -1 -1/49/10 = -1/490....?

OpenStudy (anonymous):

plz put parenthesis

OpenStudy (anonymous):

Im confused........

OpenStudy (anonymous):

I think A but not sure

OpenStudy (anonymous):

yes i got a

OpenStudy (anonymous):

Ohhh now I get it with the parentheses... -2/5 - (3/7) = -29/35 * 7/10 = (29/50)

OpenStudy (anonymous):

oh ok ty

OpenStudy (anonymous):

but thatrs not 1 of the choices

OpenStudy (anonymous):

then why would u put it on there=p

OpenStudy (anonymous):

i didnt thank you verymuch :)

OpenStudy (andijo76):

-7/10

OpenStudy (andijo76):

thats exact form -0.7 is decimal form

OpenStudy (anonymous):

oh ok ty

OpenStudy (andijo76):

welcome